OpenStudy (anonymous):

Convert 12.035 to a fraction and reduce to lowest terms (mixed number)

OpenStudy (anonymous):

i have the same question :/

OpenStudy (anonymous):

\[12.035 = \frac{12035}{1000}=\frac{2407\times5}{200\times5}=\frac{2407}{200}\]

OpenStudy (anonymous):

it has to be in mixed number form

OpenStudy (anonymous):

\[\frac{2407}{200}=12\frac{7}{200}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome